Liquid of mass 2.5 kg at temperature 20 degree celsius is heated to 40 degree Celsius find the difference intensities at this temperature if its density at 20 °C is 13.6gm/cm^3
1
Expert's answer
2020-12-08T10:53:24-0500

It is mercury.

"\Delta \rho=\rho_{20}-\rho_{40}=\rho_{20}\left(1-\frac{1}{1+\beta \Delta T}\right)\\
\Delta \rho=13.6\left(1-\frac{1}{1+(180\cdot10^{-6})20}\right)=0.0488\frac{g}{cm^3}\\
\Delta \rho=48.8\frac{kg}{m^3}"